OpenStudy (anonymous):

30 over 6 over 5 = 150 over N What is N

OpenStudy (anonymous):

\[\frac{\frac{30}{6}}{5} = \frac{150}{N}\] or \[\frac{30}{\frac{6}{5}} = \frac{150}{N}\]

OpenStudy (anonymous):

yes tht is it What is n

OpenStudy (anonymous):

which one is the correct equation that polpak posted?

OpenStudy (anonymous):

the bottom one

OpenStudy (anonymous):

Ok, so multiply both sides of the equal sign by 5 to get rid of the complex fraction. What do you get?

OpenStudy (anonymous):

Err. actually multiply the left side only by 5/5. That will rid you of the fraction in the denominator.